OpenStudy (anonymous):

I need a perfect answer for this and fast answer. Mix W-5 cups of concentrate and 8 cups of cold water. Mix X-3 cups of concentrate and 6 cups of water. Mix Y- 6 cups of concentrate and 9 cups of water. Mix Z 3 cups of concentrate and 4 cups of cold water. Question. If you make a single batch of each, what fraction of the batch would be concentrate?

OpenStudy (tkhunny):

So, \(\dfrac{W-5}{8} = \dfrac{X-3}{6} = \dfrac{Y-6}{9} = \dfrac{Z-3}{4}\) Let's see... 8 = 2*2*2 6 = 2*3 9 = 3*3 4 = 2*2 Therefore, 2*2*2*3*3 = 72 is our target. So, \(\dfrac{W-5}{8}\cdot\dfrac{9}{9} = \dfrac{X-3}{6}\cdot\dfrac{12}{12} = \dfrac{Y-6}{9}\cdot\dfrac{8}{8} = \dfrac{Z-3}{4}\cdot\dfrac{18}{18}\) Are we getting anywhere?